# The Higgs Mass and a Perspective on Broken Autonomy of Scales

### v1

@@ 3.2. The Applicability of Technical Naturalness
A third sensitivity arises already in the mesocosm itself, without the need to know anything about quantum effects: in the theory of complex systems. This represents a significant phenomenological extension of linear Newtonian physics without leaving classical physics. In linear mechanics, the dynamics of a system is unaffected by infinitesimal disturbances. The dynamics of nonlinear systems, on the other hand, show a sensitive dependence on the initial conditions with the slightest influence - small changes have a significant effect (UV-IR mixing). This is particularly evident in the butterfly effect, provocatively expressed in the question of whether the flapping of a butterfly’s wings in Brazil can trigger a tornado in Texas (Lorenz 1972). More precisely, the presence of turbulence in a multitude of complex systems, extremely dependent on tiny perturbations of initial conditions, hints that a manifestation of broken autonomy of scales occurs as a fundamental property in the mesoscopic world as well, while the linear, i.e. effective theory describes only a small fraction of deterministic phenomena that show no dependencies on distant energy scales.
Contrasted with violations of autonomy of scales, I refer to three successes of technical naturalness in quantum field theory: First, the divergent electron self-energy in classical electrodynamic, cured by the postulate of positrons by Dirac. Second, the pion mass difference, which can only be called technically natural, if the cutoff $\Lambda$ of its effective theory appears around 800 MeV, cured by the later discovered $\rho$ meson. Third, the mass difference between the long- and short-lived state of the $K$ meson (kaon). Empirical reasons demand $\Lambda<3$ GeV, which led to the prediction of the mass of the charm-quark [^10] Instead of citing all single publications, we refer to (Williams 2015), where all three phenomena are explained in greater detail.. All three successes of technical naturalness share the commonality that the phenomena to which the principle was applied must not be considered fundamental, in stark contrast to the three cases previously discussed. Instead, they each occur in effective field theories (self-explanatory for classical electrodynamics) and involve composites (mesons) of elementary particles described in a narrow interval of the energy spectrum. It is helpful to remark, as Williams (2015) has already outlined, that only the last phenomenon was explained by explicitly applying the demand for naturalness. The former ones were recognized retrospectively as confirmations of the naturalness principle in QFT.
@@
We do not know why the decoupling of distant energy scales is a widespread phenomenon that simplifies life for physicists. In addition to the mentioned cases from QFT, it holds equally true in classical physics. It is the reason why physics could gradually approach more challenging theories, and the description of an apple falling from a tree did not require quantum gravity. It is the conditio sine qua non of an effective theory. However, there is no reason to assume that (both weak and strong) autonomy of scales must be preserved at the end of the chain of effective theories. Quite the opposite: An effective theory aims to provisionally and simplifiedly represent a part of reality (a section of the entire energy scale) - a fundamental theory, on the other hand, captures elements of reality itself (not limited to specific energy ranges). A violation of autonomy of scales, in both definitions, seems almost a natural consequence [^11] It should be noted that this viewpoint is far from consensus: Crowther (2019) provides a list of criteria a fundamental theory shall fulfil and explicitly mentions naturalness as one criterion..
4. Beyond Effective Field Theories
@@
The absence of technical naturalness becomes a cause for concern only under the assumption being discussed: that the seemingly reductionist theory sequence is endless and the series of effective theories does not terminate. Lacking autonomy of scales would indeed signal the failure of reductionism. In this case, against whose plausibility we will now present arguments, the general modus operandi of foundational physics would be challenged to such an extent that it must be elevated to the status of a naturalness problem, rather than characterizing it as an almost desirable trait. We will see below how influential publications have cemented the adherence to EFTs, thereby hindering a comprehensive consideration of the hierarchy problem.
4.1. The World as a Collection of Quasi-Autonomous Domains?
@@
In the following discussion, we revisit the ontological consequences of considering autonomy of scales. To strengthen the argument for the reevaluation of the absence of technical naturalness, we turn to a widely discussed work by Cao and Schweber (1993), which describes physics as the aforementioned infinite tower of quasi-autonomous domains, each ontologically equivalent and effectively described. Thus, the authors not only adopt an anti-reductionist and anti-fundamentalist stance (infinite tower), but they also postulate an ontological pluralism, thereby inadequately representing the methods and understanding of physics in the last century, as we will see below. While in the first part of the work, they meticulously trace the history of renormalization, they derive significant consequences from Wilson’s realistic reinterpretation of formal renormalization into effective field theory (1975) in the second part. These consequences have been criticized extensively (e.g., Hartmann 2001), partly due to the use of the decoupling theorem (Appelquist, Carazzone 1975), which was excluded from the argument list for good reasons beforehand. Nevertheless, I will once again challenge the worldview outlined there, as the scope of the work has influenced the assessment of absent autonomy of scales.
It is helpful to recap the mode of operation of a foundational physicist: She deems a theory effective because she understands it as a provisional approximation to a cosmos governed by a fundamental theory. At the time of developing the EFT, this fundamental theory is evidently not within intellectual and perhaps empirical reach yet, so one settles for an approximation reliable in certain domains. During the convergence (in its empirical adequacy) of the sequence of theories reliable in certain domains against a fundamental theory reliable in all domains, it represents a precursor of merely epistemic value. Instead of making ontologically fundamental statements already, it simply propels the physicist a significant step forward on this path. This represents a central assumption of foundational physics, abandoning it would instead deny the epistemic value of such interim successes. This pragmatic approach to the world of the smallest, whose exploration is solely driven by the pure desire for knowledge and cannot be motivated by practical significance, is legitimized only by the attainability of a foundation. EFTs thus serve as instruments to approach fundamental theories. This approximation is missing in the infinitely layered tower of EFTs. The potential senselessness in the quest for fundamental physics in this picture is not an argument in itself that can be used against Cao and Schweber. It is rather the fact that ever higher precision and ever smaller distances to the extreme scales of physics (Planck scale) do not give any reason to believe that we are just as far from the foundation of the tower as in Newtonian mechanics, as suggested by the image of the infinite tower. In short, Cao and Schweber paint a false picture of the way of working of a foundational physicist in exploring the world, as well as of the world itself.
@@
Critique can also be directed at other theses: The coexistence of different EFTs by no means allows the conclusion that our world is ontologically plural. An effective theory, by definition, cannot enable ontological determinations. Moreover, no actual autonomy of these respective domains can be established. There are continuous transitions in, for example, the theory of strong nuclear force, where the approximate validity of point-like nucleons is steadily reduced, and at no discrete point does the absolute necessity of quarks become apparent. Such smooth transitions are already present in the construction of the theory to be effective. The transition to another domain is always at the discretion of the physicist, who decides what deviations are tolerable for the given context.
@@
With these arguments, the long-term significance of EFTs and thus the adherence to technical naturalness is once again questioned. Williams (2015) initially concludes that its violation in the case of the Higgs mass is a counterargument to the worldview previously presented by Cao and Schweber (1993). However, even its general consideration (without knowing anything about autonomy of scales) reveals, as shown, sufficient incoherence with the discipline of physics to reject it.
4.2. Selective Realism in the Realm of EFTs
@@
Athanasiou (2022, p. 115) categorizes the work of Cao and Schweber (1993) among that of EFT enthusiasts (among optimists, cautious and sceptics). In contrast, when addressing EFT skeptics, he states, "Clearly, the skeptical road is the path of the patient: in the absence of a completely satisfactory theory, thou shalt not make any ontological commitment". Indeed, this commandment brings us back to the initial question. Has the last word been spoken concerning the particle mass with the Higgs boson? Only in this case it was a valid argument that the absence of technical naturalness is an expression of a final, scale-crossing theory [^12] Already in 1979, Iliopoulos remarked (p. 89): ”Several people believe, and I share this view, that the Higgs scheme is a convenient parametrization of our ignorance concerning the dynamics of spontaneous symmetry breaking, and elementary scalar particles do not exist”. However, the detection of the Higgs particle is likely to refute this radical conclusion drawn from the quadratic UV sensitivity of scalar particles.. Koren (2020) concedes that a fundamental Higgs particle could indeed be the way the universe works. Nevertheless, she expects the Higgs mass to be a parameter resulting from input values of a UV theory in low-energy theory, even if it is unknown whether the Higgs mass ultimately emerges as part of a larger multiplet, as a bound state of fermions, or as an excitation in string theory. However, as we have already noted, those theories beyond the standard model, especially string theory, remain experimentally unconfirmed and pure speculation. Thus, it is hoped that the research program of BSM physics will ultimately lead to the desired situation, allowing for actual ontological statements about the smallest building blocks of matter for the first time.
@@
However, how can the presumably fundamental nature of the Higgs field in the universe be reconciled with the fact that the standard model as a whole must be assigned to the class of EFTs, making final ontological conclusions premature? An answer to this question is provided by the approach of Selective Realism, formulated by Worrall (1989) and Kitcher (1993), among others. As a subcategory of Scientific Realism, it aims to label a thoughtful selection of individual theoretical assumptions as true or at least approximately true. The argument goes as follows: Having learned from the history of science, one should be cautious of untimely euphoria and not regard the currently best theories as truthful descriptions of reality. This principle is found under the term of pessimistic meta-induction according to Laudan (1981). However, the selective realist concedes that elements can be extracted from these theories that are likely to survive theory changes unscathed and will also be a central part of (all) improved theories in the future. They are, in a sense, the fixed points in the dynamic system of changing theories. Thus, the No-Miracle Argument is taken into account, suggesting that the outstanding empirical adequacy of the theory regarding the selected aspect would be inexplicable if it cannot be attributed at least approximately true. As Musgrave points out (1988), this argument can be considered the strongest support for all varieties of realism. However, given the evident limitations of the standard model to specific energy scales (EFT), individual realism that regards the entire model as (at least approximately) true is unjustified. This does not even require the argument of pessimistic meta-induction - it was, by construction, a provisional model that will one day be discarded. Whether the future replacement of the theory as a whole is known in advance or not, a selective realist avoids Laudan’s pessimism and argues that realistically interpretable aspects were causally linked to the temporary empirical adequacy of the theory. He attributes lasting value to them for theory-building and sees them as guarantors of success for future theories.
@@
If an empirically successful aspect breaks the strong autonomy of scales, as in the case of the Higgs mass, the instrumentalist approach via EFTs, to which no even approximate truth can be ascribed, gradually comes to an end, if no experimental explanations for this behavior can be demonstrated. However, failure within the framework of such pragmatic theories need not be considered a problem, as the term "naturalness problem" has always suggested: Respecting the success of the theoretical description of the world using Higgs bosons, one should instead attribute to them a fundamental truth and cease the strained attempt to artificially force a fundamental phenomenon into the scope of provisional instrumentalist theories. Violations of strong autonomy of scales may be used as a criterion to already employ Selective Realism in the context of EFTs. However, it is crucial that the assumption of a fundamental nature of the respective phenomenon is justifiable. After all considerations, there is indeed nothing against the notion that particle masses are generated by coupling to the Higgs field on all energy scales. It is plausible that future theories will employ the same mechanism. A similar scrutiny is withstood by the previously discussed butterfly effect (an immutable principle in complex systems, also in refined future theories), noncommutative geometry (fundamental structure of spacetime in theories of quantum gravity), and the cosmological constant. On the other hand, the discussed phenomena with restored strong autonomy of scales occur in limited energetic domains (such as meson physics or classical electrodynamics). In light of this reconsideration, one can agree with Giudice (2017) that naturalness can still serve as an essential guiding principle even in the post-naturalness era, as long as its scope is understood.
@@
In contrast, for example, the fermions of the standard model can indeed be considered fundamental and irreducible, even though their quantum corrections to mass through chiral symmetries are natural. As long as there are no signs of substructures of these assumed point-like and elementary particles, Selective Realism can also be applied here [^13] To avoid misunderstandings: Of course, the OSR is a sort of Selective Realism in another sense of this term as well. The purpose of OSR in this article, however, should not be confused with the application of Selective Realism, rescueing the potential contradiction of ontological statements about objects in EFTs. The OSR was only used as a tool to determine the physical reality of bare mass and quantum corrections. This kind of Selective Realism states which kinds of entities really exist fundamentally (structures instead of fundamental objects). Selective Realism in its second meaning just states that at each point within the evolution of a theory series, a certain quotient of already realistic elements occurs in the corresponding theory, whereas the remainder can still be considered as instrumental or pragmatic, thus metaphorical elements of the theory - providing a compromise between Scientific Realism and Instrumentalism in developed, but non-final theories.. Thus, violated strong autonomy of scales is not a necessary criterion for attributing (approximately) truth to theoretical assumptions, as I would like to explicitly emphasize. However, it can be understood as a sufficient criterion once the fundamental status of the phenomenon has been sufficiently made plausible.
@@
As a perspective, we note that the lack of mathematical understanding of QFT carries the risk that its use in present-day particle physics could be fundamentally doomed. Grinbaum and Rivat (2020, p. 11) remark: "If the formalism of the new theory after a future theory change happens to be incommensurable with the mathematical structure of QFT, even the parts favored by selective realists will become void of ontological status". Even the Higgs boson with its UV-sensitive mass corrections is currently embedded in the formalism of QFT. We cannot exclude that a reformed version of QFT could shed new light on the problem of technical naturalness in its mathematical definition over $\Lambda^{2}$. The role of rigorous mathematics for an (approximately) true description of the world must therefore be thoroughly examined. Perhaps, to make ontological statements about the world, one cannot settle for effective or pragmatic mathematics. In this case, the phenomenological fundamentality of a phenomenon would not be sufficient to transition to Selective Realism. In other words, the question arises: To what extent must fundamental (non-pragmatic, not inconsistency-ignoring) mathematics and the description of fundamental phenomena go hand in hand for the theory to allow ontological statements? The background of this question is the idea that mathematics in foundational physics is not just a language for describing the world but is understood in this discipline as the world itself, an assemblage of mathematical structures. Ontology would then mean nothing less than writing down a set of mathematical equations and principles, as proposed by Tegmark (2008) among other authors. The mathematics used today to describe the (nevertheless, in our assumption, fundamental) Higgs boson, for example, could distort its true nature to such an extent that present-day physicists are not yet privy to any truth. Such distortion could occur, for example, in resolving the previously described Landau pole, which the gauge coupling to the hypercharge reaches at finite energies, prohibiting the limit $\Lambda\to\infty$. The residual jeopardy of the fundamental character of the Higgs boson in the current state of particle physics could be eliminated by the mathematics of the future, but also delay the applicability of Selective Realism.
5. Changes in the Aesthetic Canon of BSM Physics

### v2

@@ 3.2 Limited Applicability of Technical Naturalness?
  3. 1c)
    Thirdly, the specific properties of the quantum vacuum at the smallest length scales influence the expansion rate of the entire universe (vacuum energy), which undoubtedly plays out on the largest available length scales when calculating the cosmological constant (see also Craig 2022 for an overview). Unlike the balancing act performed between the electroweak scale ($10^{-18}$ m) and the Planck scale ($10^{-35}$ m) in the hierarchy problem of the Higgs mass, quantum effects in cosmology are responsible for considerably larger length scales. On the other hand, effective theories usually separate cosmology and QFT101010 This passage only addressed the principal influence of the quantum vacuum on the largest possible length scales. Moreover, the violated strong autonomy of scales exists in a way mathematically analogous to the Higgs mass: Summing the bare cosmological constant from general relativity and the value of the vacuum energy, $\tilde{\lambda}+\lambda=\Lambda$, results in handling a mismatch of 120 orders of magnitude between quantum effects and the observable value $\Lambda$ (to be distinguished from the cutoff)..
The fundamental character of the phenomena from above is easily ascribed, be it the butterfly effect (an immutable principle in complex systems, also in refined future theories), noncommutative geometry (fundamental structure of spacetime in theories of quantum gravity - "theories of everything") or the cosmological constant (determining the geometric evolution of the universe itself). Contrasted with violations of autonomy of scales in presumably fundamental phenomena, let us continue with three successful applications of the principle of technical naturalness in quantum field theory: [^11] Instead of citing all single publications, we refer to (Williams 2015), where all three phenomena are explained in greater detail. It is helpful to remark, as Williams has already outlined, that only the last phenomenon was explained by explicitly applying the demand for naturalness. The former ones were recognized retrospectively as confirmations of the naturalness principle in QFT.
  1. 2a)
    The divergent electron self-energy in classical electrodynamic, cured by the postulate of positrons by Dirac.
  2. 2b)
    The pion mass difference, which can only be called technically natural, if the cutoff $\Lambda$ of its effective theory appears around 800 MeV, cured by the later discovered $\rho$ meson.
  3. 2c)
    The mass difference between the long- and short-lived state of the $K$ meson (kaon). Empirical reasons demand $\Lambda<3$ GeV, which led to the prediction of the mass of the charm-quark
All three successes of technical naturalness share the commonality that the phenomena to which the principle was applied must not be considered fundamental, in stark contrast to the three cases previously discussed. Instead, they each occur in effective field theories (self-explanatory for classical electrodynamics) and involve composites (mesons) of elementary particles described in a narrow interval of the energy spectrum. Hence, the technical naturalness appears to be not universally applicable, but a metatheoretical guiding principle with a limited scope. From its successes we know, we derive that it can only be applied to non-fundamental phenomena modelled in provisional effective theories. This is the key message of hypothesis $H_{1}$.
@@
We do not know why the decoupling of distant energy scales is a widespread phenomenon that simplifies life for physicists. In addition to the mentioned cases from QFT, it holds equally true in classical physics. It is the reason why physics could gradually approach more challenging theories, and the description of an apple falling from a tree did not require quantum gravity. It is the conditio sine qua non of an effective theory. However, there is no reason to assume that autonomy of scales must be preserved outside of effective theories. Quite the opposite: An effective theory aims to provisionally and simplifiedly represent a part of reality (a section of the entire energy scale) - a fundamental theory, on the other hand, captures elements of reality itself (not limited to specific energy ranges). A violation of autonomy of scales, in both the weak and strong sense, seems almost a natural consequence! It must be emphasized that our viewpoint is far from consensus: Crowther (2019), for instance, provides a list of criteria a fundamental theory shall fulfil and explicitly mentions naturalness as one criterion. However, based on the comparison from above giving rise to hypothesis $H_{1}$, which arguments can be provided that the metatheoretical principle of naturalness has to be fulfilled beyond the end of the chain of EFTs - to be fulfilled in final theories describing fundamental phenomena as well? Giudice (2017, p. 8) writes "one way or another, naturalness will still play a role in the post-naturalness era". Of course, we agree: it will certainly play some role. But like other advocates of naturalness, he does not concretize the scope of applicability. This is where $H_{1}$ enters the stage and adds a first new aspect to the ongoing naturalness debate. It restricts the applicability of technical naturalness to these theories containing non-fundamental entities and relations - presumably, this criterion will play a role in future effective theories while predicting new physics. The precise formulation of this thesis is crucial - via an ontological vocabulary rather than the commonly used vocabulary of effective theories, which is closely related to it: Naturalness is of course tied to the concept of EFTs by definition. In contrast to the formulation presented here, however, the Higgs (embedded in the EFT of the standard model) inevitably raises a naturalness problem. The Higgs boson can only be clearly separated from the phenomena explained in 2a to 2c in the following chapter 4 by arguing about its status of ontological fundamentality.
Arguing for the ontologically fundamental status of the Higgs boson in the next chapter will reveal thus the hierarchy problem of the Higgs mass as a pseudo-problem, thereby deducing $H_{2}$ from $H_{1}$.
4 Placing the Higgs Outside of EFTs
@@
The absence of technical naturalness becomes a cause for concern only under the assumption discussed before and consolidated in $H_{1}$: that the seemingly reductionist theory sequence is endless and the series of effective theories does not terminate. Lacking autonomy of scales would then indeed signal the failure of reductionism. In this case, against whose plausibility we will now present arguments, the general modus operandi of foundational physics would be challenged to such an extent that missing autonomy of scales must be elevated to the status of a naturalness problem, rather than characterizing it as an almost desirable trait, possibly revealing a fundamental character of the entity of concern. We will see below how influential publications have cemented the adherence to EFTs, thereby hindering a comprehensive consideration of the hierarchy problem. But what is there to be said against the omnipresence of EFT and, over and above, for the fundamental nature of the Higgs boson?
4.1 The World is More Than a Set of Quasi-Autonomous Domains.
@@
In the following discussion, we critically evaluate the ontological consequences of considering autonomy of scales. To strengthen the argument for the reevaluation of the absence of technical naturalness, we turn to a widely discussed work by Cao and Schweber (1993), which describes physics as the aforementioned infinite tower of quasi-autonomous domains, each ontologically equivalent and effectively described. Thus, the authors not only adopt an anti-reductionist and anti-fundamentalist stance (infinite tower), but they also postulate an ontological pluralism, thereby inadequately representing the methods and understanding of physics in the last century, as we will see below. The consequences of that world view have been criticized extensively (e.g., Hartmann 2001), partly due to the use of the decoupling theorem (Appelquist, Carazzone 1975), which is today excluded in naturalness debate for good reasons beforehand. Nevertheless, we will once again challenge the worldview outlined there, as the scope of the work has influenced the assessment of absent autonomy of scales. Thereby, it seems, the dogmatic adherence to EFTs ultimately makes the hierarchy problem of the Higgs mass so persistent in the philosophical and particle physics debate. So let us liberate ourselves from the dogma that effective theories are the only way we can describe the universe.
@@
To do that, it is helpful in a first step to acknowledge that the coexistence of different EFTs by no means allows the conclusion that our world is ontologically plural. An effective theory, by definition, cannot enable ontological determinations. Moreover, no actual autonomy of these respective domains can be established. There are continuous transitions in, for example, the theory of strong nuclear force, where the approximate validity of point-like nucleons is steadily reduced, and at no discrete point does the absolute necessity of quarks become apparent. Such smooth transitions are already present in the construction of the theory to be effective. The transition to another domain is always at the discretion of the researcher, who decides what deviations are tolerable for the given context.
Let us recap the modus operandi of a foundational physicist: She deems a theory effective because she understands it as a provisional approximation to a cosmos governed by a fundamental theory. At the time of developing the EFT, this fundamental theory is evidently not within intellectual and perhaps empirical reach yet, so one settles for an approximation reliable in certain domains. During the convergence (in its empirical adequacy) of the sequence of theories reliable in certain domains against a fundamental theory reliable in all domains, it represents a precursor of merely epistemic value. Instead of making ontologically fundamental statements already, it simply propels the physicist a significant step forward on this path. This represents a central assumption of foundational physics, abandoning it would instead deny the epistemic value of such interim successes. This pragmatic approach to the microcosm, whose exploration is solely driven by the pure desire for knowledge and cannot be motivated by practical significance, is legitimized only by the attainability of a foundation. EFTs thus serve as instruments to approach fundamental theories. This approximation is missing in the infinitely layered tower of EFTs. The potential senselessness in the quest for fundamental physics in this picture is not an argument in itself that can be used against Cao and Schweber. It is rather the fact that ever higher precision and ever smaller distances to the extreme scales of physics (Planck scale) do not give any reason to believe that we are just as far from the foundation of the tower as in Newtonian mechanics, as suggested by the image of the infinite tower. In short, Cao and Schweber paint a false picture of the way of working of a foundational physicist in exploring the world, as well as of the world itself. We are certainly able to reach at least some of the fundamental elements of being, allowing for ontological commitments. Ontological plurality rules out the existence of fundamental entities, the search for which is actually one of the main tasks of fundamental physics. Pure EFTs, on the other hand (containing only instrumentalist constructions of no deeper ontological meaning), are constantly being replaced by more and more glances on the foundation of our cosmos, where the ‘tower of EFTs’ gradually ends and is replaced.
@@
With these arguments, the long-term significance of EFTs and thus the adherence to technical naturalness is once again questioned, especially questioning it as a guiding principle for "final" theories to come. Williams (2015) concludes that its violation in the case of the Higgs mass is a counterargument to the worldview previously presented by Cao and Schweber (1993). However, even its general consideration (without knowing anything about autonomy of scales) reveals, as shown, sufficient incoherence with the discipline of physics to reject it. Hence, let us move beyond effective theories and incorporate fundamental entities into them.
4.2 Selective Realism Allows for a Fundamental Higgs in Preliminary EFTs
@@
Athanasiou (2022, p. 115) categorizes the work of Cao and Schweber (1993) into that of EFT-enthusiasts (among optimists, cautious and sceptics). In contrast, when addressing EFT-sceptics, he states, "Clearly, the sceptical road is the path of the patient: in the absence of a completely satisfactory theory, thou shalt not make any ontological commitment". Indeed, this commandment brings us back to the initial question: Has the last word already been spoken concerning the real, the true nature of the Higgs boson? Only in that case we had a valid argument that the absence of technical naturalness in its renormalized mass can be interpreted as a sign of having reached final, scale-crossing theory without causing any "naturalness problem" (the essence of $H_{2}$, mainly deduced from $H_{1}$) - as for noncommutative spacetime or senstive dependence on initial conditions of compley systems (cf. Section 3.2). Koren (2020) expects the Higgs mass to be a parameter resulting from input values of a UV theory in low-energy theory, even if it is unknown whether the Higgs mass ultimately emerges as part of a larger multiplet, as a bound state of fermions, or as an excitation in string theory. However, as we have already noted, those theories beyond the standard model, especially string theory, remain experimentally unconfirmed and pure speculation. There is no convincing reason to believe that the Higgs boson might not be a ontologically fundamental entity. Thus, it is hoped that the research program of BSM physics will ultimately lead to the desired situation, allowing for actual ontological commitments about the smallest building blocks of matter for the first time.
@@
However, how can the presumably fundamental nature of the Higgs field in the universe - the crucial assumption for $H_{2}$ be reconciled with the fact that the standard model as a whole must be assigned to the class of EFTs, making final ontological conclusions premature? The answer to this last question, which ultimately enables our envisaged diagnosis of a pseudo-problem, will be provided in the following by the approach of Selective Realism, formulated by Worrall (1989) and Kitcher (1993), among others. As a subcategory of Scientific Realism, it aims to label a thoughtful selection of individual theoretical assumptions as true or at least approximately true, while others are still preliminary, provisional constructions. Hence, Selective Realism claims that at least some entities contained in a still not final theory must be already ontologically fundamental (or simply: real). The argument goes as follows: Having learned from the history of science, one should be cautious of untimely euphoria and not regard the currently best theories as truthful descriptions of reality. This principle is found under the term of pessimistic meta-induction according to Laudan (1981). However, the selective realist concedes that elements can be extracted from these theories that are likely to survive theory changes unscathed and will also be a central part of (all) improved theories in the future. They are, in a sense, the fixed points in the dynamic system of changing theories. Thus, the No-Miracle Argument is taken into account, suggesting that the outstanding empirical adequacy of the theory regarding the selected aspect would be inexplicable if it cannot be attributed at least approximately true. As Musgrave points out (1988), this argument can be considered the strongest support for all varieties of realism. However, given the evident limitations of the standard model to specific energy scales (EFT), classical Scientific Realism that regards the entire model as (at least approximately) true is unjustified. This does not even require the argument of pessimistic meta-induction - it was, by construction, a provisional model that will one day be discarded. Whether the future replacement of the theory as a whole is known in advance or not, a selective realist avoids Laudan’s pessimism and argues that realistically interpretable aspects were causally linked to the temporary empirical adequacy of the theory. He attributes lasting value to them for theory-building and sees them as guarantors of success for future theories.
@@
If an empirically successful element of a theory breaks autonomy of scales (as in the case of the Higgs mass), the instrumentalist approach via EFTs (as the standard model as a whole), to which no even approximate truth can be ascribed, gradually comes to an end, if no experimental explanations for this behavior can be demonstrated. However, failure within the framework of such pragmatic theories need not be considered a problem, as the term "naturalness problem" has always suggested: Keeping in mind the "natural" occurrence of UV-IR-mixtures in the examples fom Section 3.2 ($H_{1}$) and simultaneously respecting th success of the theoretical description of the world using Higgs bosons up to very high energy scales (no-miracle-argument), one should instead attribute to the Higgs boson a fundamental truth (i.e., a realistic interpretation) and cease the dogmatic attempt to artificially force a fundamental phenomenon into the scope of purely provisional, instrumentalist theories. This is ultimately the justification of $H_{2}$. Violations of strong autonomy of scales may thus be used as an encouragement to replace the purely instrumental EFT approaches by Selective Realism. In this view, it sends a positive signal to all the physicists searching for fundamental entities. After all considerations, there is indeed nothing against the notion that particle masses are generated by coupling to the Higgs boson on all energy scales. It is highly plausible instead, that future theories will employ the same mechanism turning the Higgs boson into an entity that will survive any theory change to come - in the spirit of Selective Realism.
@@
One last, but crucial distinction shall be noted: The fermions of the standard model can still be considered fundamental and irreducible, even though their quantum corrections to their mass are natural through chiral symmetries (’t Hooft natural). As long as there are no signs of substructures of these assumed point-like and elementary particles, Selective Realism can also be applied here [^12] To avoid misunderstandings: Of course, the OSR is a sort of Selective Realism in another sense of this term as well. The purpose of OSR in this article, however, should not be confused with the application of Selective Realism, rescueing the potential contradiction of ontological statements about objects in EFTs. The OSR was only used as a tool to determine the physical reality of bare mass and quantum corrections. This kind of Selective Realism states which kinds of entities really exist fundamentally (structures instead of fundamental objects). Selective Realism in its second meaning just states that at each point within the evolution of a theory series, a certain quotient of already realistic elements occurs in the corresponding theory, whereas the remainder can still be considered as instrumental or pragmatic, thus metaphorical elements of the theory - providing a compromise between Scientific Realism and Instrumentalism in developed, but non-final theories.. Thus, violated strong autonomy of scales is not a necessary criterion for attributing (approximate) truth to theoretical assumptions. It must be rather understood as a sufficient criterion once the fundamental nature of the phenomenon has been sufficiently made plausible.
@@
As a perspective and challenge to the application of Selective Realism to QFT in general, we note that the lack of mathematical understanding of QFT carries the risk that its use in present-day particle physics could be fundamentally doomed. Grinbaum and Rivat (2020, p. 11) remark: "If the formalism of the new theory after a future theory change happens to be incommensurable with the mathematical structure of QFT, even the parts favored by selective realists will become void of ontological status". Even the Higgs boson with its UV-sensitive mass corrections is currently embedded in the formalism of QFT. We cannot exclude that a reformed version of QFT could shed new light on the problem of technical naturalness in its mathematical definition over $\Lambda^{2}$. The role of rigorous mathematics for an (approximately) true description of the world must therefore be thoroughly examined [^13] Perhaps, to make ontological statements about the world, one cannot settle for effective or pragmatic mathematics. In this case, the phenomenological fundamentality of a phenomenon would not be sufficient to transition to Selective Realism. In other words, the question arises: To what extent must fundamental (non-pragmatic, not inconsistency-ignoring) mathematics and the description of fundamental phenomena go hand in hand for the theory to allow ontological statements? The background of this question is the idea that mathematics in foundational physics is not just a language for describing the world but is understood in this discipline as the world itself, an assemblage of mathematical structures. Ontology would then mean nothing less than writing down a set of mathematical equations and principles, as proposed by Tegmark (2008) among other authors. The mathematics used today to describe the (nevertheless, in our assumption, fundamental) Higgs boson, for example, could distort its true nature to such an extent that present-day physicists are not yet privy to any truth. Such distortion could occur, for example, in resolving the previously described Landau pole, which the gauge coupling to the hypercharge reaches at finite energies, prohibiting the limit $\Lambda\to\infty$. The residual jeopardy of the fundamental character of the Higgs boson in the current state of particle physics could be eliminated by the mathematics of the future, but also delay the applicability of Selective Realism..
5 Conclusion
This article provided novel (metatheoretical) arguments that the restoration of technical naturalness of the Higgs mass with new physics represents an engagement with a pseudo-problem. However, reinterpreting the failure of the guiding principle of technical naturalness in case of the Higgs boson allowed for ontological commitments about high-energy physics: a statement about its status as a fundamental entity of our cosmos - a realist element embedded into an effective field theory (the standardmodel) being by definition an instrumentalist construction. Major justifaction came from incorporating the concept of Selective Realism into the standard model, claiming that there are already fundamental (real) entities and relations that will survive any future theory change, included into that overall effective framework. Hence, the smallness of the Higgs mass reveals itself as a glimmer of hope: A hint at the fundamental nature of the affected phenomenon, which to reveal is the long-term mission of fundamental physics. We believe that as long as one only dealt with the artificial restoration of strong autonomy of scales in the corrections to the Higgs mass through theories without empirical evidence, there is a degenerating element in the BSM research program. This article denies the status of a problem while simultaneously enabling a progressive shift. In this shift, the limit of the applicability of pure EFTs is accepted, legitimizing greater openness in theory-building and evaluation without attacking the general approach of BSM physics through effective theories: What impact does the discussion about the gradual dissolving (purely) effective theories have on the near future of BSM physics in general? Regarding the forthcoming use of EFTs, we can only align ourself with the pragmatism of Georgi (1993, p. 215) writing: "In addition to being a great convenience, effective field theory allows us to ask all the really scientific questions that we want to ask without committing ourselves to a picture of what happens at arbitrarily high energy". Indeed, the broken autonomy of scales in the correction to the Higgs mass should be understood as a failure of the adherence to pure EFTs for this specific scenario. Selective Realism provides a compromise without being overly optimistic about the persistence of the standard model as a whole. The more generall use of EFTs remains a central prerequisite for a thriving particle physics in the future. This article merely criticized the dogmatic adherence to principles for which there are convincing reasons to release them: The simple and, given the current empirical evidence, plausible assumption of a Higgs boson that is ontologically fundamental, suffices (with $H_{1}$ and $H_{2}$) to turn the hierarchy problem into a pseudo-problem.
Appendix A: Changes in the Aesthetic Canon of BSM Physics